# ToolFlow: Boosting LLM Tool-Calling Through Natural and Coherent Dialogue Synthesis

### v1

@@ Abstract
Zezhong Wang1thanks: Work done during the internship at Huawei Noah’s Ark Lab., Xingshan Zeng2, Weiwen Liu2, Liangyou Li2, Yasheng Wang2, Lifeng Shang2, Xin Jiang2, Qun Liu2, Kam-Fai Wong1 1The Chinese University of Hong Kong, 2Huawei Noah’s Ark Lab {zzwang,kfwong}@se.cuhk.edu.hk {zeng.xingshan,liuweiwen8,liliangyou}@huawei.com {wangyasheng,Shang.Lifeng,Jiang.Xin,qun.liu}@huawei.com
@@ 5.1 Settings
Baselines Simple Multiple Parallel Parallel multiple Avg.
Claude-3.5-Sonnet 88.55 95.00 91.50 92.50 91.22
GPT-4-turbo-0409 87.45 96.50 91.00 89.00 90.28
GPT-4o-0513 80.55 91.00 90.00 83.00 85.02
LLaMA-3.1 8B 90.36 89.50 73.50 73.50 83.40
[…]
@@ 5.1.1 Datasets
  • BFCL Patil et al. (2023) primarily consists of Python-style tool call data, divided into four categories Simple, Multiple, Parallel, and Parallel Multiple. We selected the categories that can be evaluated with the Abstract Syntax Tree (AST), which are statistically stable and easy to evaluate. The accuracy is reported.
@@ 5.2 Results
5.2.1 ToolFlow achieves tool-calling capability comparable to GPT-4.
@@
The results are shown in Tabel 5.1. Compared to the baseline, ToolFlow achieved performance comparable to most mainstream LLMs. ToolFlow slightly outperformed GPT-4 and GPT-4o, but was slightly weaker than Claude-3.5-Sonnet. The difference is mainly reflected in Parallel Multiple type questions. In these types of questions, the model is presented with multiple candidate tools and required to call these tools in parallel, which is of high complexity in nature. We believe that the difference in the model sizes results in ToolFlow’s performance being inferior to Claude, as ToolFlow only has 8B parameters.
The ablation experiment shows that the model trained on data synthesized by strategies including both Graph-based Sampling and Generated Plan performs the best. This is especially evident in Parallel Multiple type questions. Compared to LLaMA-3.1-8B-Instruct, the trained model achieved an accuracy improvement of 11.5% on Parallel Multiple (increasing from 73.5% to 85%).
@@ A.1 Details of Data Quality Assessment
A.2 Prompts and Demonstrations
See Table 8 to 12 for details.
A.3 Train with other base model with ToolFlow
Table 13 and 14 display the results of two other base models fine-tuned with data generated with ToolFlow.
Table 8: Prompt for plan generation
Table 9: Prompt for GPT-4 Data Evaluation
Table 10: Prompt for User agent
Table 11: Prompt for Assistant agent
Table 12: Prompt for Tool agent

### v2

@@ Abstract
Zezhong Wang1thanks: Work done during internship at Huawei Noah’s Ark Lab., Xingshan Zeng2thanks: Corresponding author, Weiwen Liu2, Liangyou Li2, Yasheng Wang2, Lifeng Shang2, Xin Jiang2, Qun Liu2, Kam-Fai Wong1 1The Chinese University of Hong Kong, 2Huawei Noah’s Ark Lab {zzwang,kfwong}@se.cuhk.edu.hk {zeng.xingshan,liuweiwen8,liliangyou}@huawei.com {wangyasheng,Shang.Lifeng,Jiang.Xin,qun.liu}@huawei.com
@@ 4.2 Quality Evaluation
4.3 Comparison with Natural Dialogue Dataset
To better understand how our synthetic dialogues compare with human-created ones, we conducted a comparative study with an established dataset. We chose the MultiWOZ dataset Budzianowski et al. (2018) as a natural dialogue dataset for comparison. MultiWOZ (Multi-Domain Wizard-of-Oz) is a well-known task-oriented dialogue dataset, and we believe comparisons with this dataset would be convincing. We repeated the GPT-4 evaluation experiment by first randomly sampling 200 dialogues from MultiWOZ. Then, using the same scoring prompt, we had GPT-4 evaluate these dialogues across the four dimensions. The results are shown in Table 3.
MultiWOZ scores slightly higher on naturalness, coherence, and accuracy compared to the ToolFlow dataset, though the differences are minimal (average score differences between 0.1-0.2). Regarding helpfulness scores, ToolFlow outperformed MultiWOZ by 0.3 points. These results suggest that our synthetic dialogues in ToolFlow achieve comparable quality to human-created dialogues in MultiWOZ, with particularly strong performance in task-oriented aspects such as helpfulness.
Dataset NAT COH HELP ACC
ToolFlow 3.72 3.91 4.71 4.92
MultiWOZ 3.98 4.03 4.41 4.95
Table 3: Evaluation Scores Comparison between ToolFlow and MultiWOZ.
@@ 5.1.1 Datasets
  • BFCL-v2 Patil et al. (2023) primarily consists of Python-style tool call data, divided into four categories Simple, Multiple, Parallel, and Parallel Multiple. Version 2 adds more data from dynamic, real world scenarios. We selected the categories that can be evaluated with the Abstract Syntax Tree (AST), which are statistically stable and easy to evaluate. The accuracy is reported.
@@ 5.2 Results
5.2.1 ToolFlow achieves tool-calling capability comparable to GPT-4o.
@@
The results are shown in Tabel 5.1. Overall, ToolFlow achieves performance comparable to GPT-4o. On the Non-Live subset, ToolFlow outperformed GPT-4 and GPT-4o, but was slightly weaker than Claude-3.5-Sonnet. On the Live subset, ToolFlow still lags behind these leading closed-source LLMs. This is because the Live subset added more user-contributed test cases from the real world, thus making it more challenging for the model. We attribute this gap primarily to differences in model size, given that ToolFlow only has 8B parameters. The ablation experiment shows that the model trained on data synthesized by strategies including both Graph-based Sampling and Generated Plan performs the best. This is especially evident in Parallel Multiple type questions.
@@ 6 Correlation Analysis
7 Dataset Overlap Analysis
To ensure the reliability of the evaluation, we conducted an overlap analysis between training and test datasets. This examination helps verify the independence of these test data and prevents potential data leakage issues. We employed both N-gram-based and similarity-based methods to demonstrate that there is no significant data leakage in the ToolFlow dataset. We also included the well-known xLam agent training set Zhang et al. (2024) as a control group for comparison.
N-gram-based method Following the approach used in LLaMA-2 Touvron et al. (2023), we considered a token contaminated if it appeared in any token n-gram longer than 10 tokens in both the evaluation sample and the training set. A tool was classified as leaked if more than 10% of the tokens in its JSON string were contaminated.
Similarity-based method We defined a tool as leaked if the cosine similarity between the given tool and any tool in the evaluation dataset exceeded 0.9. We used the all-MiniLM-L12-v2 encoder from HuggingFace [^2] https://huggingface.co/ to obtain representations for all tools.
We present the proportions of data leakage across different evaluation metrics in Table 9. These results suggest that there is no severe data leakage between ToolFlow as a training set and the test sets.
Training ToolFlow xLam Test N-gram Similarity N-gram Similarity BFCL 0.239% 2.922% 0.656% 5.247% APIBank 0.000% 1.887% 0.000% 3.774% ToolAlpaca 0.000% 0.000% 0.000% 0.000%
Table 9: Overlap between Training and Test Sets
@@ 8 Conclusion
Acknowledgements
This work was partially supported by Hong Kong RGC GRF No. 14206324, CUHK direct grant No. 4055209, and CUHK Knowledge Transfer Project Fund No. KPF23GWP20.
@@ A.1 Details of Data Quality Assessment
A.2 Reliability Analysis of GPT-4 as an Evaluator
To validate the reliability of GPT-4’s evaluations, we conducted additional human evaluations. Specifically, we sampled 50 examples from the data previously evaluated by GPT-4 for human assessment. We recruited four Computer Science PhD students to rate these 50 samples using the same criteria as GPT-4: naturalness, coherence, helpfulness, and accuracy. We collected the human evaluation results and calculated both the Cohen’s Kappa inter-rater agreement scores among evaluators and the Pearson correlation coefficients between human and GPT-4 evaluations.
Metric NAT COH HELP ACC
Cohen’s Kappa 0.63 0.71 0.92 0.95
Pearson Correlation 0.61 0.76 0.84 0.86
Table 10: Correlation between Human and GPT-4 Ratings
The results in Table 10 demonstrate high inter-rater agreement and strong correlations with GPT-4’s evaluations, particularly for helpfulness and accuracy metrics. While the agreement and correlations for naturalness and coherence were relatively lower compared to the other two metrics, they still maintained a minimum correlation of 0.61, indicating a strong positive correlation between human and GPT-4 evaluations. We believe these results substantiate the reliability of GPT-4’s evaluations.
A.3 GPT-4 v.s. Open Source LLM for Data Synthesis
To evaluate the reliance of synthesis on closed-source, high-performance LLMs (e.g., GPT-4), we conducted comparative experiments using open-source alternatives. Following the same pipeline but replacing GPT-4 with LLaMA-3.1-8B-instruct, we conducted preliminary synthesis experiments. While still in the initial validation phase, we have synthesized 4,235 dialogue instances. We replicated experiments using this data to fine-tune LLaMA-3.1-8B-instruct, with BFCL test results shown in Table 11.
Data Source Simple Multiple Parallel Parallel multiple Avg.
N/A 90.36 89.50 73.50 73.50 83.40
LLaMA-3.1-8B-instruct 89.88 90.10 85.55 80.15 86.42
GPT-4 91.23 91.85 87.10 84.45 88.66
Table 11: Performance comparison of models trained on data synthesized by GPT-4 and LLaMA.
@@ A.2 Reliability Analysis of GPT-4 as an Evaluator
Here, "N/A" represents the baseline performance of LLaMA-3.1-8B-instruct (consistent with results reported in the main text). Notably, we leveraged existing results from Section 6’s Correlation analysis, where we had downsampled to ten subsets of 4,000 dialogues each, comparable to our LLaMA-3.1-8B-instruct synthesized dataset size. The table reports the average results across these ten subsets. While data synthesized by LLaMA-3.1-8B-instruct indeed shows lower performance compared to GPT-4-synthesized data, the gap is not substantial. Moreover, compared to LLaMA-3.1-8B-instruct’s initial performance, training on its self-synthesized data demonstrates improvements in tool-calling capabilities.
A.4 Train with other base model with ToolFlow
Table 12 and 13 display the results of two other base models fine-tuned with data generated with ToolFlow. Results demonstrate that ToolFlow enhances tool-calling performance across multiple LLMs, validating the generalizability of our synthetic data approach.
A.5 Prompts and Demonstrations
See Table 14 to 18 for details.
Table 14: Prompt for plan generation
Table 15: Prompt for GPT-4 Data Evaluation
Table 16: Prompt for User agent
Table 17: Prompt for Assistant agent
Table 18: Prompt for Tool agent